OpenStudy (anonymous):

sqrt(x)=11sqr(y) Find dy/dx? Can someone give me the answer to this?

OpenStudy (kropot72):

\[\sqrt{x}=11\times \sqrt{y}\] Squaring both sides we get \[x=121y\] \[y=\frac{x}{121}\] Can you find dy/dx now?